There are 1/4 red grapes + 3/8 green grapes + 2/16 blue grapes + 16 blue grapes
How many in total
The answer given was 64
The common denominator used was 64.
16 red grapes, 24 green grapes, 8 blue grapes + 16 blue grapes
With fractions, shouldn't you find the least common denominator. In this case 16.
Thanks

Answer by Boreal(15235)

Total grapes = x
(1/4)x+(3/8)x+(1/8)x+16=x
(3/4)x+16=x
16=(1/4)x
x=64
Having the blue grapes twice is an issue because if you end up with 4 red 6 green and 2 blue using 16, you still have another 16 blue to add in.
Fractions have to be evaluated as what they mean.
How many grapes are present is the question.
Then you're given fractions of that unknown quantity. Then you solve. Otherwise, you are just adding up fractions without knowing what you have. It isn't a problem of 1/4 +2/3 +...
it is a problem of 1/4 of the total +2/3 of the Total + ... that is a very different problem.